Find the probability of selecting the same card from a deck of playing cards twice with replacement.

Find the probability of selecting the same card from a deck of playing cards twice without replacement.

Answer by rothauserc(4718)

with replacement
:
Probability(selecting same card twice) = 1/52 * 1/52 = 1/2704
:
without replacement
:
Probability(selecting same card twice) = 1/52 * 0 = 0
: